On a map two cities are 5 1/2 inches apart.The map uses a scale of 2 inches = 75 miles.

What is actual distance between two cities?

Answer by Theo(13342)

i believe you mean 2 inches = 75 miles.

if so, then you get the ratio 2 inches / 75 miles = 5 and 1/2 inches / x

that ratio looks like 2 / 75 = 5.5 / x

cross multiply to get 2 * x = 75 * 5.5

solve for x to get x = 75 * 5.5 / 2 = 206.25 miles.

another way to look at it is.

the map uses a scale of 2 inches = 75 miles.

divide both sides of that equation by 2 and you get 1 inch = 37.5 miles.

in order words, you get 37.5 miles in the real world to 1 inch on the map.

multiply 37.5 * 5.5 and you get 206.25 miles.